[figure]

N_{p} N_{q}

N_{x} N_{s} N_{l} N_{t}

T T S

2^{(N_{c}N_{x}N_{t})^{a}} a ≥ 1 n_{c} n_{c} ∝

N_{c}^{3}N_{x}^{2}N_{t}^{3} ,

N_{c} e_{1} e_{2} e_{1} e_{2} e_{1} e_{2} ≥

0

G = (V, E)

S

C′ C

C′ ⊆ C

C′ C′

Using summation is not strictly necessary, provided that the components can be combined in some fashion to reflect the quality of the solution. N_{s} s k = s s C(s)

T_{0} T = T_{0}

s′ ⊂ N_{s} Δ =

C(s′)− C(s)

Δ ≤ 0 {

} s =

s′ Δ > 0 {

} s =

s′ =

e^{−Δ/(kT)} T s